# Q-Net: Transferable Queue Length Estimation via Kalman-based Neural Networks

### v1

@@ 7 Conclusion
In this study, we proposed Q-Net, a computationally and data-efficient framework for queue length estimation. Our problem setting reflects a general and practically relevant scenario in which traffic flow is only partially observed. Moreover, Q-Net relies solely on two widely available and privacy-preserving data sources: cumulative vehicle counts from loop detectors and aggregated floating car data (aFCD). Together, these aspects highlight the broad applicability of the proposed method.
The filtering algorithm of Q-Net leverages KalmanNet, an AI-augmented Kalman filter that combines the representational power of deep learning with the structure of state-space modeling. Unlike traditional Kalman filters, KalmanNet does not require prior knowledge of system or noise dynamics; instead, it learns to adapt dynamically to real-world noise patterns. Our contributions extend beyond a direct application of KalmanNet in three key ways: (a) Q-Net introduces a tailored state-space formulation specifically designed for queue length estimation. This formulation effectively fuses heterogeneous data sources with differing spatial and temporal resolutions. (b) Q-Net adapts KalmanNet to decouple measurement dimensionality from segment length, thereby enabling spatial transferability across different road segments. (c) Q-Net employs a training strategy designed to handle long time series data, ensuring stable and efficient model convergence.
Q-Net was implemented and evaluated on multiple road sections in Rotterdam, the Netherlands. Empirical results show that it consistently outperforms baseline methods during both the formation and dissipation phases of traffic queues, effectively correcting the lag typically observed in aFCD-based estimates. The model demonstrated strong temporal and spatial transferability, underscoring its practicality for real-world deployment without the need to maintain or install costly radar infrastructure. Furthermore, a real-time variant of Q-Net was introduced, illustrating its potential for dynamic traffic control applications.
Future research could enhance Q-Net by integrating graph machine learning to capture spatial correlations in road networks with multiple intersections, enabling more expressive evolution models for improved queue change estimation. Integrating additional contextual features, such as traffic signal timings or weather conditions, may further boost accuracy. Moreover, exploring alternative training paradigms, such as curriculum learning, and validating the approach across a more diverse set of urban environments would strengthen generalizability. Ultimately, combining Q-Net with traffic control and optimization strategies opens new avenues for intelligent, data-driven traffic management systems.

### v2

@@ 8 Conclusion
This paper addresses the challenge of estimating queue length in partially observed environments by proposing Q-Net, a Kalman-based framework that integrates model-driven and data-driven principles. Evaluated on multiple main road sections in Rotterdam, Q-Net consistently outperforms baseline methods and effectively corrects the lag typically observed in aFCD-based estimates. The model’s strong performance across temporal, spatial, and real-time online settings validates our hybrid approach. These results demonstrate Q-Net’s practicality as a robust solution for real-world deployment, offering accurate queue estimation without the need to maintain or install costly radar infrastructure.
Acknowledgments
The authors gratefully acknowledge the support of the Rotterdm municipality for providing the data. This research has been produced by the “EMERALDS” project which has received funding under the Horizon Europe research and innovation programme (GA No. 101093051).
Appendix A Data split
Table 6: All data were collected in November 2023. The table presents the specific days assigned to the training, validation, and test sets in each split.
Split Train days Validation days Test days
1 16, 17, 10, 27, 23, 13, 21, 15, 24, 25, 19 20, 11 14, 22, 18
2 14, 22, 16, 17, 15, 24, 20, 21, 10, 18, 11 23, 25 27, 13, 19
3 14, 24, 13, 23, 15, 17, 20, 10, 27, 19, 11 22, 18 21, 16, 25
4 27, 20, 21, 23, 14, 17, 10, 22, 16, 19, 18 13, 25 15, 24, 11
5 20, 10, 22, 14, 16, 27, 15, 21, 13, 19, 11 24, 18 23, 17, 25
6 27, 22, 23, 24, 21, 15, 16, 17, 14, 11, 25 13, 19 18, 10, 20